# Evolution Meets Diffusion: Efficient Neural Architecture Generation

### v1

@@ 4 Experiments
In Section 4.1, we conduct sufficient experiments in NAS-BENCH-101 and NAS-BENCH-201 search spaces and extend our study to TransNASBench-101-Micro/Macro and MobileNetV3 search space. Subsequently, we perform ablation experiments on the FD strategy, as detailed in Section 4.2. Finally, further experiments on generation efficiency and neural predictors are carried out in Section 4.3.
@@ 4.1.1 Experiment Setup
For the generation process, the denoising step of EDNAG is 100. The noise scale of diffusion ranges from $\alpha_{T}=10^{-4}$ to $\alpha_{0}=1-10^{-4}$ and the noise scale of denoising is $\sigma=0.8$. The population size (or batch size) is 100 for MobileNetV3 search space and 30 for other search spaces.
4.1.2 Experiments on NAS-Bench-101 Search Space
Table 1: Experiments on NAS-BENCH-101 search space. Metrics are reported over 3 runs on CIFAR-10 with 95% confidence intervals.
Method Val Acc Test Acc
RS Bergstra and Bengio (2012) 93.92$\pm$0.11 93.42$\pm$0.02
RL Williams (1992) 94.23$\pm$0.33 93.72$\pm$0.27
RE Real et al. (2019) 94.28$\pm$0.24 93.78$\pm$0.16
FADZS-NAS Wang et al. (2024) 93.80$\pm$0.08 93.22$\pm$0.08
DiNAS Asthana et al. (2024) 94.83$\pm$0.03 94.40$\pm$0.07
EDNAG (Ours) 94.86$\pm$0.00 94.46$\pm$0.01
Global Best 94.86 94.47
NAS-Bench-101 is a typical search space encompassing $4.23\times 10^{5}$ unique neural architectures. It provides up to 7 layers with 3 candidate operations and a maximum of nine inter-layer connections.
We conduct experiments on NAS-Bench-101 for image classification task of CIFAR-10, displayed in Table 1. EDNAG outperforms previous methods, particularly surpassing the prior state-of-the-art approach (DiNAS), by improving validation and test accuracy by 0.03% and 0.06% respectively, which highlights the strong capability of EDNAG in generating task-optimal architectures.
@@ 4.1.3 Experiments on NAS-Bench-201 Search Space
NAS-BENCH-201 is a widely used cell-based search space comprising 15,625 distinct neural architectures. Each cell in NAS-Bench-201 includes 6 layers and 5 predefined operations (zeroize, skip-connect, 1×1 conv, 3×3 conv, 3×3 avg pool) for each layer.
We perform conditional generation of architectures across four diverse task datasets within the NAS-BENCH-201 search space, as shown in Table 2. EDNAG generates neural architectures and subsequently trains the top-performing architecture on the task datasets to validate their performance. In experiments on CIFAR-10 and CIFAR-100, EDNAG achieves comparable performance on test accuracies compared to previous baselines. On Aircraft and Pets, EDNAG surpasses previous baselines with improvements of 2.23% and 10.45% on test accuracies, respectively. Across diverse task datasets, EDNAG outperforms all previous NAS and NAG methods, achieving state-of-the-art (SOTA) performance and showcasing exceptional capability in generating optimal architectures. This is attributed to the FD strategy, which efficiently guides architecture samples toward high-fitness subspaces during the iterative generation process, serving as a powerful conditional mechanism to steer architectures toward global optimality. Compared to DiffusionNAG, the previous SOTA baseline that directly generates neural architectures using naive diffusion models with limited optimization abilities, EDNAG excels with its superior global optimization capabilities while maintaining efficient generative abilities.
Moreover, EDNAG also significantly improves the efficiency of architecture validation by reducing the number of architectures (Archs) requiring training to achieve final accuracy. For CIFAR-10 and CIFAR-100, the final accuracies are obtained from the NAS-Bench-201 benchmarks Dong and Yang (2020) by querying only the top-1 generated architecture. For Aircraft and Pets, only the top-1 architecture requires full training in EDNAG, enhancing validation efficiency by 3$\times$ and 2$\times$ compared to DiffusionNAG. Results illustrate that the top-1 neural architecture identified by our neural predictor attains optimal accuracy after training and testing, demonstrating its remarkable performance, achieved by the application of coreset selection prior to dataset encoding. Unlike the random sampling method used in DiffusionNAG, this approach significantly enhances the stability and reliability of the predicted accuracies.
@@ 4.1.4 Experiments on TransNASBench-101 Search Space
Table 3: Experiments on TransNASBench-101 Search Space across seven tasks. Metrics are reported with 95% confidence intervals over 3 runs. For improved readability, the values in the table for SSIM, mIoU, and L2 Loss are scaled by 100.
@@
TransNASBench-101 extends downstream tasks beyond basic image classification, encompassing a diverse set of 7 tasks: Object Classification, Scene Classification, Autoencoding, Surface Normal Estimation, Semantic Segmentation, Room Layout Prediction, and Jigsaw Puzzle Solving. Moreover, it defines two distinct types of search spaces: a cell-based micro search space and a stack-based macro skeleton search space, which encompass 4,096 and 3,256 unique architectures, respectively.
We conduct experiments in both micro-level and macro-level search spaces, as presented in Table 3. Across both search spaces, EDNAG exhibits significant performance improvements compared to previous baselines and consistently achieves globally optimal results across all downstream tasks while maintaining remarkable stability across three independent runs. These findings not only demonstrate the effectiveness of the FD strategy in improving global optimization capabilities for generative models, but also highlight the transferability of EDNAG for cross-task neural architecture generation within various downstream scenarios.
@@ 4.1.5 Experiments on MobileNetV3 Search Space
Due to the relatively limited scale of the above search spaces, our study further extends to the larger and more complex MobileNetV3 search space to demonstrate the scalability of EDNAG. MobileNetV3 is a large search space which encompasses over $2.17\times 10^{19}$ distinct neural architectures, making it more than $10^{15}$ times larger than the NAS-Bench-201 search space. MobileNetV3 offers five stages, each consisting of 2 to 4 building blocks, where each building block is defined by the kernel size of convolutional layers chosen from {3, 5, 7} and the expansion rate of output channels chosen from {3, 4, 6}.
We conduct experiments on four distinct task datasets within the MobileNetV3 search space, as shown in Table 4. In CIFAR-10, CIFAR-100, Aircraft, and Pets, EDNAG demonstrates an improvement in accuracy of up to 0.49%. This experiment not only demonstrates the superior ability of architecture generation for EDNAG but also confirms its adaptability and effectiveness in large-scale search spaces. This is attributed to the fact that when transitioning to different search spaces, DiffusionNAG necessitates retraining the score network within the diffusion model to accommodate the noise matrix of the new dimensionality, which is troublesome and time-consuming. In contrast, as a training-free NAG method, EDNAG avoids the need for training or retraining, thereby enabling adaptation to new search spaces with negligible costs.
@@ 4.2 Ablation Study
Naive diffusion models often results in convergence to local optima, resulting in suboptimal architecture generation. The integration of the FD strategy significantly enhances the accurancy of generated architectures, facilitating the generation of globally optimal solutions. In addition, the incorporation of the optimization strategy substantially accelerates the convergence of the denoising process. These results underscore the effectiveness of both the FD strategy and the optimization strategy within EDNAG, demonstrating their pivotal roles in achieving task-optimal architecture generation.
@@ 4.3.1 Comparison of Generation Efficiency
As the first network-free neural architecture generation method, EDNAG boasts remarkable characteristics, including minimal computational resource consumption and rapid conditional generation, etc. To demonstrate its efficiency, we compare EDNAG with DiffusionNAG in terms of computational costs during both training and inference processes in the NAS-Bench-201 search space, as shown in Table 5. Since EDNAG is free of training, it takes 0 GPU hours for the training process, while DiffusionNAG requires 3.43 GPU hours in the NAS-BENCH-201 search space on a Tesla V100-SXM2. For generation process, we evaluate the inference time of both EDNAG and DiffusionNAG on an NVIDIA RTX 4090. EDNAG achieves speedups of up to 100$\times$, 100$\times$, 7$\times$, and 5$\times$ compared to DiffusionNAG on CIFAR-10, CIFAR-100, Aircraft and Pets, respectively. Therefore, EDNAG not only enables neural architecture generation without time-consuming training processes but also significantly reduces computational cost during the inference process. This is attributed to the FD strategy during the generative process, which directs the denoising process without relying on a score network, thereby achieving remarkable execution efficiency. In contrast, DiffusionNAG depends on a complex score network at each step of the generative process to predict the noise to be removed, which demands extensive and frequent network inference, significantly slowing down the generation process.
@@ 4.3.2 Quality Analysis of Generated Architectures
We generate 1000 neural architectures by EDNAG and DiffusionNAG, then we display their distribution in Figure 6. Compared to DiffusionNAG, EDNAG demonstrates a higher proportion of high-performing architectures. This demonstrates that the architectures generated by EDNAG exhibit a denser distribution within high-accuracy subspaces, reflecting a marked improvement in the quality of architecture generation. This advancement is achieved through the powerful conditional guidance and superior optimization ability by combining evolutionary algorithms, which contrasts sharply with the straightforward use of naive diffusion models in DiffusionNAG.
@@ 4.3.3 Analysis of Neural Predictor
Figure 7: Performance analysis of neural predictor. We predict the accuracy of architectures for CIFAR-10 and CIFAR-100, and display the Pearson correlation coefficient between predicted and actual accuracies.
Figure 8: Stability analysis of neural predictor. With neural predictors in DiffusionNAG and EDNAG, we assessed the accuracy of identical neural architectures on consistent task datasets across 100 repeated runs.
We conduct experiments within the NAS-BENCH-201 search space and report Pearson correlation coefficients between predicted and actual accuracies, as illustrated in Figure 7. On CIFAR-10, the predictor of DiffusionNAG suffers from biased predictions due to random sampling, resulting in a lower correlation coefficient, while EDNAG achieves a 12.09% improvement. On CIFAR-100, the predictor of EDNAG achieves a 7.29% enhancement compared to DiffusionNAG. This experiment reveals that the neural predictor of EDNAG, which utilizes K-Means clustering to select coreset images for dataset encoding, aligns more closely with real architectural accuracy. This approach effectively circumvents the limitations of random sampling in DiffusionNAG, which fails to preserve the uniformity of the target dataset distribution, resulting in biased predictions.
Moreover, during the denoising iterations, identical architectures in successive generations ($x_{t}$ and $x_{t-1}$) necessitate multiple evaluations of the same architecture. Our experiments, shown in Figure 8, evaluate the same neural architecture over 100 repeated runs. The results demonstrate significant fluctuations in the predicted accuracy of DiffusionNAG due to random image sampling. In contrast, EDNAG demonstrates consistent predicted accuracy, effectively avoiding random perturbations in fitness values and further enhancing the stability of the whole generation process.
@@ 5 Conclusion
This study introduces an Evolutionary Diffusion-based Neural Architecture Generation Approach, called EDNAG, an efficient neural architecture generation method without networks. We propose a fitness-guided denoising strategy (FD) that maps fitness to probabilistic space, thereby guiding the DDIM-based denoising process for architecture generation. Extensive experiments demonstrate that EDNAG achieves SOTA performance in generating task-optimal architectures and surpasses existing baselines in generation efficiency and computational cost. We believe that EDNAG will significantly contribute to the development of NAS by reducing computational costs and accelerate its practical application.

### v3

@@ 4 Experiments
In Section 4.1, we conduct sufficient experiments in NAS-BENCH-101, NAS-BENCH-201, and NAS-Bench-301 search spaces and extend our study to TransNASBench-101-Micro/Macro and MobileNetV3 search spaces. Subsequently, we perform ablation experiments on the FD strategy, as detailed in Section 4.2. Finally, further experiments on generation efficiency and neural predictors are carried out in Section 4.3. For sensitivity analysis of parameters, comprehensive comparisons of GPU time across 14 baselines, and more experiments, please refer to the Appendix.
@@ 4.1.1 Experiment Setup
For the generation process, the denoising step of EDNAG is 100. The noise scale of diffusion ranges from $\alpha_{T}=10^{-4}$ to $\alpha_{0}=1-10^{-4}$ and the noise scale of denoising is $\sigma=0.8$. The population size (or batch size) is 100 for MobileNetV3 search space and 30 for other search spaces. More implementation details, search spaces description, and introduction to baselines can be found in the Appendix.
@@ 4.1.2 Experiments on NAS-Bench-201 Search Space
NAS-BENCH-201 is a widely used cell-based search space comprising 15,625 distinct neural architectures. Each cell in NAS-Bench-201 includes 6 layers and 5 predefined operations for each layer. Within the NAS-BENCH-201 search spacee, we perform conditional generation of architectures across four task datasets and subsequently train the top-performing architecture to validate their performance, as shown in Table 1. In experiments on CIFAR-10 and CIFAR-100, EDNAG achieves comparable performance on test accuracies compared to previous methods. In Aircraft and Pets, EDNAG surpasses previous baselines with improvements of 2.23% and 10.45% on test accuracies, respectively. Across diverse task datasets, EDNAG outperforms all previous NAS and NAG methods, achieving state-of-the-art (SOTA) performance and showcasing exceptional capability in generating optimal architectures. This is attributed to the FD strategy, which efficiently guides architecture samples toward high-fitness subspaces during the iterative generation process, serving as a powerful conditional mechanism to steer architectures toward global optimality. Compared to DiffusionNAG, the previous SOTA baseline that directly generates neural architectures using naive diffusion models with limited optimization abilities, EDNAG excels with its superior global optimization capabilities while maintaining efficient generative abilities.
Moreover, EDNAG also significantly improves the efficiency of architecture validation by reducing the number of architectures (Archs) requiring training to achieve final accuracy. For CIFAR-10 and CIFAR-100, the final accuracies are obtained from the NAS-Bench-201 benchmarks Dong and Yang (2020) by querying only the top-1 generated architecture. For Aircraft and Pets, only the top-1 architecture requires full training in EDNAG, enhancing the validation efficiency of 3$\times$ and 2$\times$ compared to DiffusionNAG. The results illustrate that the top-1 neural architecture identified by our neural predictor attains optimal accuracy after training and testing, demonstrating its remarkable performance, achieved by the application of coreset selection before dataset encoding. Unlike the random sampling method used in DiffusionNAG, this approach significantly enhances the stability and reliability of the predicted accuracies.
4.1.3 Experiments on NAS-Bench-101 and NAS-Bench-301 Search Space
Table 2: Experiments on NAS-BENCH-101 and NAS-Bench-301 search space. Accuracies (%) are reported over 3 runs on CIFAR-10 with 95% confidence intervals.
Method NAS-Bench-101 NAS-Bench-301
RS Bergstra and Bengio (2012) 93.59$\pm$0.42 94.40$\pm$0.08
AG-Net Xiao and Wang (2024) 94.18$\pm$0.07 94.45$\pm$0.12
SRLNAS Cassimon et al. (2025) 93.86$\pm$0.14 94.53$\pm$0.04
DiNAS Asthana et al. (2024) 94.27$\pm$0.20 94.54$\pm$0.15
EDNAG (Ours) 94.46$\pm$0.01 94.64$\pm$0.05
Global Best 94.47 94.69
NAS-Bench-101 is a cell-based search space encompassing $4.23\times 10^{5}$ architectures while NAS-Bench-301 is a DARTS Liu et al. (2018)-like search space consisting over $10^{10}$ architectures.
We conduct experiments on NAS-Bench-101 and NAS-Bench-301 for image classification task on CIFAR-10, displayed in Table 2. EDNAG outperforms previous methods, particularly surpassing previous SOTA methods, by improving accuracy by 0.2% and 0.1% respectively, which highlights the strong capability of EDNAG in generating task-optimal architectures.
@@ 4.1.4 Experiments on TransNASBench-101 Search Space
Table 3: Experiments on TransNASBench-101 across seven tasks. Metrics are reported with 95% confidence intervals over 3 runs. For improved readability, the values in the table for SSIM, mIoU, and L2 Loss are scaled by 100.
@@
TransNASBench-101 extends downstream tasks beyond basic image classification, encompassing a diverse set of 7 tasks: Object Classification, Scene Classification, Autoencoding, Surface Normal Estimation, Semantic Segmentation, Room Layout Prediction, and Jigsaw Puzzle Solving. Moreover, it defines two types of search spaces: a cell-based micro search space and a stack-based macro skeleton search space, which encompass 4,096 and 3,256 unique architectures, respectively.
We conduct experiments in both micro-level and macro-level search spaces, as presented in Table 3. Across both search spaces, EDNAG demonstrates substantial performance enhancements of up to 7.02% over prior baselines, consistently approaching global optimality across all downstream tasks while maintaining stability across three independent runs. These findings not only demonstrate the effectiveness of the FD strategy in improving global optimization capabilities for generative models, but also highlight the transferability of EDNAG for cross-task neural architecture generation within various downstream scenarios.
@@ 4.1.5 Experiments on MobileNetV3 Search Space
Due to the relatively limited scale of the above search spaces, our study further extends to a larger and more complex MobileNetV3 search space to demonstrate the scalability of EDNAG. MobileNetV3 is a large search space which encompasses over $2.17\times 10^{19}$ distinct neural architectures, making it more than $10^{15}$ times larger than the NAS-Bench-201 search space.
We conduct experiments on four distinct task datasets within the MobileNetV3 search space, as shown in Table 4, where EDNAG demonstrates an improvement in accuracy of up to 0.49%. This experiment not only demonstrates the superior ability of architecture generation for EDNAG but also confirms its adaptability and effectiveness in large-scale search spaces. This adaptability is attributed to the fact that when transitioning to different search spaces, DiffusionNAG necessitates retraining the score network within the diffusion model to accommodate the new dimensionality of noise matrix, which is troublesome and time-consuming. In contrast, as a training-free NAG method, EDNAG avoids the need for training or retraining, thereby enabling adaptation to new search spaces with negligible costs.
@@ 4.2 Ablation Study
Naive diffusion models often result in convergence to local optima, resulting in suboptimal architecture generation. The integration of the FD strategy significantly enhances the accurancy of generated architectures, facilitating the generation of globally optimal solutions. In addition, the incorporation of the optimization strategy substantially accelerates the convergence of the denoising process. These results underscore the effectiveness of both the FD strategy and the optimization strategy within EDNAG, demonstrating their pivotal roles in achieving task-optimal architecture generation.
@@ 4.3.1 Comparison of Generation Efficiency
As the pioneering network-free NAG approach, EDNAG achieves minimal computational resource consumption and rapid conditional generation. To demonstrate its efficiency, we compare EDNAG with DiffusionNAG in terms of computational costs during both training and inference processes in the NAS-Bench-201 search space, as shown in Table 5. Since EDNAG is free of training, it takes 0 GPU hours for the training process, while DiffusionNAG requires 3.43 GPU hours in the NAS-BENCH-201 search space on a Tesla V100-SXM2. For generation process, we evaluate the inference time of both EDNAG and DiffusionNAG on an NVIDIA RTX 4090. EDNAG achieves speedups of up to 100$\times$, 100$\times$, 7$\times$, and 5$\times$ compared to DiffusionNAG on CIFAR-10, CIFAR-100, Aircraft and Pets, respectively. Therefore, EDNAG not only enables neural architecture generation without time-consuming training processes but also significantly reduces computational cost during the inference process. This is attributed to the FD strategy during the generation process, which directs the denoising process without relying on a score network, thereby achieving remarkable execution efficiency. In contrast, DiffusionNAG depends on a complex score network at each step of the generative process to predict the noise to be removed, which demands extensive and frequent network inference, significantly slowing down the generation process.
@@ 4.3.2 Quality Analysis of Generated Architectures
We generate 1000 neural architectures by EDNAG and DiffusionNAG, then display their distribution in Figure 5. Compared to DiffusionNAG, EDNAG demonstrates a higher proportion of high-performing architectures. This shows that the architectures generated by EDNAG exhibit a denser distribution within high-accuracy subspaces, reflecting a marked improvement in the quality of architecture generation. This advancement is achieved through the powerful conditional guidance and superior optimization ability by combining evolutionary algorithms, which guides architecture samples into a more stable and optimal subspace.
@@ 4.3.3 Analysis of Neural Predictor
Figure 6: Pearson correlation coefficients analysis of neural predictor.
Figure 7: Stability analysis of neural predictor over 100 repeated runs.
We report Pearson correlation coefficients between predicted and actual accuracies in NAS-BENCH-201, as illustrated in Figure 7. On CIFAR-10 and CIFAR-100, the predictor of DiffusionNAG suffers from biased predictions due to random sampling, resulting in a lower correlation coefficient, while EDNAG achieves improvements of 12.09% and 7.29%, respectively, by K-Means clustering to select coreset images before dataset encoding.
Moreover, during denoising iterations, identical architectures in successive generations ($x_{t}$ and $x_{t-1}$) require multiple evaluations of the same architecture. Our experiments in Figure 7, evaluate a same neural architecture over 100 repeated runs. The results show significant fluctuations in the predicted accuracy of DiffusionNAG due to random image sampling. In contrast, EDNAG demonstrates nearly consistent predicted accuracy with its variance reduced by 48$\times$, effectively avoiding random perturbations in fitness values and further enhancing the stability of the whole generation process.
@@ 5 Conclusion
This study introduces an Evolutionary Diffusion-based Neural Architecture Generation Approach, called EDNAG, an efficient neural architecture generation method without networks. We propose a fitness-guided denoising strategy (FD) that maps fitness to probabilistic space, thereby guiding the DDIM-based denoising process for architecture generation. Extensive experiments demonstrate that EDNAG achieves SOTA performance in generating task-optimal architectures and surpasses existing baselines in generation efficiency and computational cost. We believe that EDNAG will contribute significantly to the development of NAS by reducing computational costs and accelerating practical application.